OpenStudy (anonymous):

seven less than x is the same as the difference of 4 more than twice x and x^2. what is x

OpenStudy (anonymous):

is this right 7-x+4-2x-x^2=x

OpenStudy (anonymous):

seven less than x means \[x-7\]

OpenStudy (anonymous):

my teacher said it doesn't matter what way you put it x-7 is the same as 7-x

OpenStudy (anonymous):

then your teacher is on the moon

OpenStudy (anonymous):

i want to go to the moon please

OpenStudy (anonymous):

let me try again k

OpenStudy (anonymous):

x-7=(4+2x)-x^2

OpenStudy (anonymous):

if your teacher thinks that x - 7 is the same as 7 - x, then your teacher thinks 10 - 7=7 - 10 very odd indeed

OpenStudy (anonymous):

she said they equal when you take magnitude of both sides

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

well that is true. but that hardly means that -3 = 3 does it?

OpenStudy (anonymous):

ok so how do i solve this

OpenStudy (anonymous):

\[x-7=(4+2x)-x^2\]

OpenStudy (anonymous):

is it x^2+x-7-(4+2x)=0

OpenStudy (anonymous):

x^2-x-11=0

OpenStudy (anonymous):

lets get rid of the parentheses first because they are unnecessary

OpenStudy (anonymous):

how do i solve this using the Way of the Myininaya?

OpenStudy (anonymous):

\[x-7=4+2x-x^2\] \[x^2-x-11=0\]

OpenStudy (anonymous):

my friend told me about it and im interested

OpenStudy (anonymous):

that is what you have yes? i would use quadratic formula for this one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i can do it thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

x=(1 + sqrt{45})/2 x=(1 -sqrt{45}}/2 i think thats right